OpenStudy (anonymous):

if \[\log_{9}4=a\]and \[\log_{27}5=b\] find an expression in terms of a & b for \[\log_{3}80\]

OpenStudy (anonymous):

i had that backwards \[2\log_9(4)=\log_3(4)\] and \[3\log_{27}(5)=\log_3(5)\] now lets see what we can do with \(\log_3(80)\)

OpenStudy (anonymous):

since \(80=4^2\times 5\) we can write \[\log_3(80)=\log_3(4^2\times 5)=\log_3(4^2)+\log_3(5)=2\log_3(4)+\log_3(5)\]

OpenStudy (anonymous):

now substitute back for \(a\) and \(b\) and you will get your answer

OpenStudy (anonymous):

this is how they explained for me to do it...so im sort of just going by this. lookin at what you did... im gettng a little lost... /: